# Matryoshka: Learning to Drive Black-Box LLMs with LLMs

### v1

@@ 4.3 Reasoning: GSM8K
Methods ($\downarrow$) gpt-3.5 4o-mini gpt-3.5 4o-mini
@@
Matryoshka 0.911 0.956 0.738 0.779
w/o IGO 0.896 0.954 0.729 0.780
Table 3: Accuracy (%) on the mathematical reasoning task using the GSM8K dataset.
@@ 4.4 Planning: ALFWorld
Main Results. Matryoshka consistently outperforms existing baselines, achieving state-of-the-art performance with an overall success rate of 95.52% on ALFWorld tasks (Table 4). This superior performance indicates that Matryoshka effectively generates plans to guide the task execution of the black-box model, enhancing its ability to interact with the environment. Furthermore, we observe that Matryoshka exhibits superior performance compared to both the untuned white-box model (w/o Guidance Optimization) and the white-box models trained with fewer rounds of Intermediate Guidance Optimization (w/o $1^{\text{st}}/2^{\text{nd}}$-round IGO). As the number of IGO training rounds increases, Matryoshka’s performance on ALFWorld correspondingly improves, ultimately raising the success rate from 81.34% to 95.52%. These results underscore the efficacy of the IGO training employed in Matryoshka.
BUTLER (Shridhar et al., 2020) 46.00 39.00 74.00 100.00 22.00 24.00 37.00
ReAct (Yao et al., 2023) 37.50 64.52 69.57 42.86 38.89 17.65 47.76
Reflexion (Shinn et al., 2023) 50.00 41.94 65.22 52.38 66.67 47.06 52.99
AdaPlanner (Sun et al., 2024a) 100.00 93.55 78.26 95.24 66.67 88.24 88.06
Matryoshka 100.00 93.55 100.00 90.48 100.00 88.24 95.52
w/o $2^{\text{nd}}$-round IGO 100.00 93.55 100.00 100.00 83.33 88.24 94.78
w/o $1^{\text{st}},2^{\text{nd}}$-round IGO 100.00 93.55 86.96 95.24 55.56 88.24 88.06
@@
[…]
Ablation Studies on Sample Efficiency. Figure 6(a) illustrates the relationship between success rate (%) and the proportion of training data used to optimize the controller. In the ALFWorld environment, Matryoshka achieves an accuracy of 94.78% using only one-quarter of the training data, surpassing the best-performing baseline, AdaPlanner, by 6.7%. This demonstrates the sample efficiency of Matryoshka in achieving high performance with limited training data. This study demonstrates that Matryoshka significantly reduces the reliance on high-quality task planning samples and expert trajectories, thereby providing a more resource-efficient solution.
@@
Figure 6: Success rate (%) w.r.t number of (a) training samples and (b) interaction turns.
Effect of Number of Interaction Turns $M$. We evaluate the performance of Matryoshka against the vanilla LLaMA3-8B-Instruct and AdaPlanner on ALFWorld, varying the number of closed-loop iterations $M$ during the inference phase. As illustrated in Figure 6(b), following DPO training, Matryoshka achieves an accuracy exceeding 95% in the open-loop inference setting ($M$=1), significantly surpassing both AdaPlanner and LLaMA3-8B-Instruct. Furthermore, during an 8-iteration closed-loop inference, Matryoshka maintains the highest accuracy of 97%. These findings indicate that Matryoshka is capable of generating exceptionally high-quality plans, enabling the GPT model serving as the task executor to interact with the environment and complete tasks successfully without requiring closed-loop refinement.
@@ 5 Related Works
Black-Box LLMs Generation Enhancement. Existing approaches aiming to enhance the generation capabilities of black-box LLMs can be broadly categorized into two groups: (1) ICL- and (2) adapter-based methods. ICL-based methods (Sun et al., 2024a; Tan et al., 2024a; Zhuang et al., 2024b) are designed to augment the original query with carefully crafted instructions or meticulously constructed few-shot demonstrations to guide the model. While this enables the black-box LLM to exhibit specific capabilities or adhere to particular directives, these methods require significant human effort in prompt engineering and result in prompts that are rigid and static. Adapter-based methods (Sun et al., 2024b; Shi et al., 2024a; Zhuang et al., 2024b) follow a best-of-N selection evaluation paradigm (Lightman et al., 2023). Given a problem, adapter-based methods generate $N$ candidate solutions from the generator and subsequently evaluate them using a lightweight adapter to identify the highest-scoring solution as the final answer. However, such methods are heavily dependent on the generative capabilities of the black-box LLM, which may result in selecting a suboptimal candidate as the best of a bad bunch.
Small LMs Drive LLMs Generation. SuperICL (Xu et al., 2023) incorporates outputs from smaller language models (LMs) as complementary information for input queries, integrating them into the context provided to black-box LLMs. However, these smaller LMs are fixed and can only support classification tasks that rely on label predictions with associated confidence scores. HYDRA (Zhuang et al., 2024b) is a retrieval-augmented generation framework that trains a BERT-sized reranker to reorder retrieved passages to better cater to user-specific requirements. Nevertheless, these methods apply only discrete optimization on the prompt through reranking and selection of few-shot demonstrations, which limits the potential improvements achievable via prompt engineering.
@@
Reinforcement Learning for Prompt Optimization. As LLMs scale, new capabilities emerge, enabling models to learn tasks efficiently through a few in-context demonstrations. To harness these capabilities, several approaches have been proposed to leverage reinforcement learning for improved prompt generation, enhancing LLM performance. RLPrompt (Deng et al., 2022) introduces an RL-based framework for generating optimal prompts via black-box optimization. Similarly, TEMPERA (Zhang et al., 2023) formulates prompt optimization as test-time prompt editing, using RL to efficiently explore the editing space. BDPL (Diao et al., 2023) further advances this by proposing a variance-reduced policy gradient algorithm to estimate gradients of parameters in the categorical distribution of each discrete prompt. However, these methods primarily focus on classification tasks, where gradient estimation is straightforward, limiting their applicability to more complex generation tasks requiring long-horizon solutions.
6 Conclusion and Future Work
@@
We introduced Matryoshka, a lightweight white-box LLM controller designed to augment the capabilities of large-scale black-box LLMs across a wide range of complex tasks, including reasoning, planning, and personalization. By leveraging a controller-generator framework with environmental feedback, Matryoshka effectively decomposes complex tasks and guides black-box LLMs through intermediate guidance. Through policy gradient optimization, Matryoshka exhibits a self-improving nature that continually enhances LLM capabilities via multi-turn guidance optimization. Extensive experiments on three diverse datasets demonstrate its effectiveness in steering black-box LLMs for long-horizon tasks without requiring access to model parameters or output probabilities. Compared to the best-performing state-of-the-art baselines, Matryoshka achieves average improvements of 3.19% in reasoning tasks, 7.46% in planning tasks, and 5.82% in personalization tasks. These results underscore the potential Matryoshka as a transparent and scalable solution, enabling white-box LLMs to drive black-box LLMs in complex problem-solving. Future work could extend Matryoshka to tackle more complex applications requiring long-horizon generation and reasoning, such as solving software engineering problems and proving mathematical theorems. Additionally, the controller component of Matryoshka could be developed into a self-enhancing mechanism or a universal controller applicable to a wide range of real-world applications.
7 Reproducibility Statement
The datasets utilized in this study are all publicly available, including LaMP for the personalization task, GSM8K and GSM-Hard for the reasoning task, and ALFWorld for the planning task. Detailed descriptions of these datasets and their corresponding tasks are provided in Appendix C. In Appendix D, we outline the baselines used for comparison and describe the experimental setup. Appendix E offers an in-depth explanation of the main experiments, including hardware and software configurations, hyperparameter settings, and step-by-step procedures for the three tasks. Additionally, Appendix F presents case studies for each task, demonstrating the superior performance of our method compared to the baselines in a more intuitive manner. Appendix G details the prompts used for each task. The implementation of Matryoshka is provided in the supplementary materials and will be released publicly available on GitHub upon acceptance.
8 Ethics Statement
@@
We strictly followed the data usage guidelines for interactions with Microsoft Azure’s OpenAI API and Gemini API service. Although our research relied solely on publicly available datasets, we took extra precautions to minimize any potential risk of information leakage. Specifically, we opted out of the human review process by completing and submitting the Azure OpenAI Additional Use Case Form [^2] https://aka.ms/oai/additionalusecase. This proactive measure highlights our commitment to maintaining the highest data privacy standards and ethical research practices, especially concerning personalization tasks.
Appendix A Limitations and Broader Impacts
A.1 Limitations
In this study, we propose a modular framework, Matryoshka, that leverages a lightweight white-box LLM controller to enhance the capabilities of black-box LLMs. Despite its effectiveness, we have identified several potential limitations of Matryoshka:
Malign Usage. Since Matryoshka employs a white-box LLM controller to augment black-box LLMs, there are notable risks to consider. Malicious actors could exploit this approach to engineer harmful capabilities or generate toxic content for training purposes. While black-box LLMs are designed to resist producing such content, our controller could be misused to manipulate these models into generating undesirable outputs. Furthermore, there is a risk that the intermediate guidance produced by our controller could be exploited to extract sensitive information from black-box LLMs, potentially facilitating jailbreaking or other targeted attacks.
Data Privacy. Matryoshka preserves the confidentiality of training data by avoiding third-party API sharing, thereby safeguarding the integrity of training samples during the enhancement process of black-box LLMs. However, when applied to personalization tasks, it is important to recognize that retrieved historical records or the queries themselves may inadvertently contain sensitive information, potentially risking unintended disclosure of private data.
A.2 Broader Impacts
@@
Potential Positive Societal Impacts. The proposed Matryoshka framework addresses a critical challenge in consistently enhancing the capabilities of black-box LLMs for long-horizon tasks with broad scopes. By improving reasoning, planning, and personalization, Matryoshka can deliver significant benefits across various domains. For instance, it can provide insights into complex theorems, advance industrial automation, and offer more personalized interactions for end users. Overall, Matryoshka has the potential to facilitate more useful, relevant, and satisfying interactions, thereby improving productivity, decision-making, and quality of life. Moreover, Matryoshka operates without requiring access to the model weights of black-box LLMs, making the technology accessible to a wide range of off-the-shelf LLM APIs and enabling seamless integration into diverse use cases. By leveraging existing LLMs, Matryoshka can be readily adopted by researchers, developers, and organizations, accelerating the development and deployment of advanced language models in real-world applications.
@@
Appendix B Additional Related Works
RLHF. Proximal policy optimization (PPO) (Schulman et al., 2017) is the predominant deep reinforcement learning method used in RLHF, leading to significant successes in models like InstructGPT (Ouyang et al., 2022), ChatGPT (Achiam et al., 2023), and Gemini (Reid et al., 2024). However, applying PPO requires extensive effort and resources (Choshen et al., 2019; Engstrom et al., 2020; Tang et al., 2024), often beyond the scope of open-source capabilities. To simplify implementation and streamline the training process, recent works (Azar et al., 2024; Ethayarajh et al., 2024) have proposed direct preference learning algorithms following the DPO framework (Rafailov et al., 2024b). These algorithms bypass the reward modeling step and directly optimize carefully designed loss objectives on the preference dataset, hence the term direct preference learning.

### v2

@@ 4.3 Reasoning: GSM8K
Table 3: Accuracy on the mathematical reasoning task using the GSM8K dataset.
@@
Method ($\downarrow$) gpt-3.5 4o-mini gpt-3.5 4o-mini
@@
M-Pilot 0.931 0.964 0.761 0.801
w/o IDPO 0.896 0.954 0.729 0.780
@@
This improvement stems from the optimized intermediate guidance generated by M-Pilot. Conditioned on this guidance, M-Pilot enables the black-box LLM to generate long-horizon solutions to solve the tasks. Similar to LaMP, M-Pilot trained with gpt-3.5-turbo can be seamlessly applied to other black-box models for solving mathematical problems on GSM8K without additional training costs. Notably, M-Pilot learns high-level planning abilities without focusing on specific details.
@@ 4.4 Planning: ALFWorld
Table 4: Success rate ($\%$) across six planning tasks from AlfWorld. For all baselines, including M-Pilot, we utilize gpt-3.5-turbo as the black-box LLM.
Methods ($\downarrow$) Tasks ($\rightarrow$) Pick Clean Heat Cool Exam Pick2 All
BUTLER (shridhar2020alfworld, ) 46.00 39.00 74.00 100.00 22.00 24.00 37.00
ReAct (yao2023react, ) 37.50 64.52 69.57 42.86 38.89 17.65 47.76
Reflexion (shinn2023reflexion, ) 50.00 41.94 65.22 52.38 66.67 47.06 52.99
AdaPlanner (sun2024adaplanner, ) 100.00 93.55 78.26 95.24 66.67 88.24 88.06
M-Pilot 100.00 93.55 100.00 95.24 100.00 88.24 96.27
w/o $2^{\text{nd}}$-round IDPO 100.00 93.55 100.00 100.00 83.33 88.24 94.78
w/o $1^{\text{st}},2^{\text{nd}}$-round IDPO 100.00 93.55 86.96 95.24 55.56 88.24 88.06
@@
Main Results. M-Pilot consistently outperforms existing baselines, achieving state-of-the-art performance with an overall success rate of 96.27% on ALFWorld tasks (Table 4). This superior performance indicates that M-Pilot effectively generates plans to guide the task execution of the black-box model, enhancing its ability to interact with the environment. Furthermore, we observe that M-Pilot exhibits superior performance compared to both the untuned white-box model (w/o Guidance Optimization) and the white-box models trained with fewer rounds of Iterative Direct Preference Optimization (w/o $1^{\text{st}}/2^{\text{nd}}$-round IDPO). As the number of IDPO training rounds increases, M-Pilot’s performance on ALFWorld correspondingly improves, ultimately raising the success rate from 81.34% to 96.27%. These results underscore the efficacy of the IDPO in M-Pilot.
@@
Figure 5: Success rate (%) w.r.t number of (a) training samples and (b) inner-loop interaction turns.
Ablation Studies on Sample Efficiency. Figure 5(a) illustrates the relationship between success rate (%) and the proportion of training data used to optimize the controller. In the ALFWorld environment, M-Pilot achieves an accuracy of 94.78% using only one-quarter of the training data, surpassing the best-performing baseline, AdaPlanner, by 6.7%. This demonstrates the sample efficiency of M-Pilot in achieving high performance with limited training data. This study demonstrates that M-Pilot significantly reduces the reliance on high-quality task planning samples and expert trajectories, improving the resource efficiency.
As illustrated in Figure 5(b), following DPO training, M-Pilot achieves an accuracy exceeding 95% in the open-loop inference setting (inner loop MM=1), significantly surpassing both AdaPlanner and LLaMA3-8B-Instruct. Furthermore, during an 8-iteration closed-loop inference, M-Pilot maintains the highest accuracy of 97%. These findings indicate that M-Pilot is capable of generating exceptionally high-quality plans, enabling the GPT model serving as the executor to interact with the environment and complete tasks successfully without closed-loop refinement.
@@ 5 Related Works
Black-Box LLMs Generation Enhancement. Existing approaches aiming to enhance the generation capabilities of black-box LLMs can be broadly categorized into two groups: (1) ICL- and (2) adapter-based methods. ICL-based methods (sun2024adaplanner, ; tan2024democratizing, ; zhuang2024hydra, ) are designed to augment the original query with well-crafted instructions or well-constructed few-shot demonstrations to guide the model. While this enables the black-box LLM to exhibit specific capabilities, these methods require significant human effort in prompt engineering and result in prompts that are rigid and static. Adapter-based methods (sunbbox, ; shi2024medadapter, ; zhuang2024hydra, ) follow a best-of-N selection evaluation paradigm (lightman2023let, ), which evaluate $N$ candidate solutions with a lightweight adapter and identify the highest-scoring solution as the final answer. However, such methods are heavily dependent on the generative capabilities of the black-box LLM, which may result in selecting a suboptimal candidate as the best of a bad bunch.
@@
Reinforcement Learning for Prompt Optimization. As LLMs scale, new capabilities emerge, enabling models to learn tasks efficiently through a few in-context demonstrations. To harness these capabilities, several approaches have been proposed to leverage reinforcement learning for improved prompt generation, enhancing LLM performance. RLPrompt (deng2022rlprompt, ) introduces an RL-based framework for generating optimal prompts via black-box optimization. Similarly, TEMPERA (zhang2023tempera, ) formulates prompt optimization as test-time prompt editing, using RL to efficiently explore the editing space. BDPL (diao2023blackbox, ) further advances this by proposing a variance-reduced policy gradient algorithm to estimate gradients of parameters in the categorical distribution of each discrete prompt. However, these methods primarily focus on classification tasks, where gradient estimation is straightforward, limiting their applicability to more complex generation tasks requiring long-horizon solutions.
6 Conclusion
@@
We introduced Matryoshka Pilot (M-Pilot), a lightweight white-box LLM controller designed to augment the capabilities of large-scale black-box LLMs across a wide range of complex tasks, including reasoning, planning, and personalization. By leveraging a controller-generator framework with environmental feedback, M-Pilot effectively decomposes complex tasks and guides black-box LLMs through intermediate guidance. Through policy gradient optimization, M-Pilot exhibits a self-improving nature that continually enhances LLM capabilities via multi-turn guidance optimization. Extensive experiments on three diverse datasets demonstrate its effectiveness in steering black-box LLMs for long-horizon tasks without requiring access to model parameters or output probabilities. Compared to the best-performing state-of-the-art baselines, M-Pilot achieves average improvements of 3.19% in reasoning tasks, 7.46% in planning tasks, and 5.82% in personalization tasks. These results underscore the potential M-Pilot as a transparent and scalable solution, enabling white-box LLMs to drive black-box LLMs in complex problem-solving.
Appendix A Impact Statement
A.1 Ethical Aspects
@@
We strictly followed the data usage guidelines for interactions with Gemini API and ChatGPT API service. Although our research relied solely on publicly available datasets, we took extra precautions to minimize any potential risk of information leakage. Specifically, we opted out of the human review process by completing and submitting the Additional Use Case Form [^2] https://aka.ms/oai/additionalusecase. This proactive measure highlights our commitment to maintaining the highest data privacy standards and ethical research practices, especially concerning personalization tasks.
A.2 Future Societal Consequences
@@
Potential Positive Societal Impacts. The proposed M-Pilot framework addresses a critical challenge in consistently enhancing the capabilities of black-box LLMs for long-horizon tasks with broad scopes. By improving reasoning, planning, and personalization, M-Pilot can deliver significant benefits across various domains. For instance, it can provide insights into complex theorems, advance industrial automation, and offer more personalized interactions for end users. Overall, M-Pilot has the potential to facilitate more useful, relevant, and satisfying interactions, thereby improving productivity, decision-making, and quality of life. Moreover, M-Pilot operates without requiring access to the model weights of black-box LLMs, making the technology accessible to a wide range of off-the-shelf LLM APIs and enabling seamless integration into diverse use cases. By leveraging existing LLMs, M-Pilot can be readily adopted by researchers, developers, and organizations, accelerating the development and deployment of advanced language models in real-world applications.
@@
A.3 Limitations
In this study, we propose a modular framework, M-Pilot, that leverages a lightweight white-box LLM controller to enhance the capabilities of black-box LLMs. Despite its effectiveness, we have identified several potential limitations of M-Pilot:
Malign Usage. Since M-Pilot employs a white-box LLM controller to augment black-box LLMs, there are notable risks to consider. Malicious actors could exploit this approach to engineer harmful capabilities or generate toxic content for training purposes. While black-box LLMs are designed to resist producing such content, our controller could be misused to manipulate these models into generating undesirable outputs. Furthermore, there is a risk that the intermediate guidance produced by our controller could be exploited to extract sensitive information from black-box LLMs, potentially facilitating jailbreaking or other targeted attacks.
Data Privacy. M-Pilot preserves the confidentiality of training data by avoiding third-party API sharing, thereby safeguarding the integrity of training samples during the enhancement process of black-box LLMs. However, when applied to personalization tasks, it is important to recognize that retrieved historical records or the queries themselves may inadvertently contain sensitive information, potentially risking unintended disclosure of private data.
Appendix B Additional Experiments
B.1 Further Ablation Studies on LaMP
Figure 6: Effect of # history per user in LaMP-2M and -4.
To further investigate the effect of user profile count on the generation of intermediate outputs, we analyze performance across different numbers of profiles per user. Figure 6 presents the accuracy and ROUGE-L curves separately for LaMP-2M and LaMP-4, with the x-axis representing the total number of profiles per user (e.g., “0-20” indicates users with 0 to 20 profiles). We compared the results of our proposed method, M-Pilot, and PAG, utilizing the white-box controller Llama-3-8B-Instruct and the black-box model gpt-4o-mini.
On LaMP-2M, as the profile count increases, PAG’s performance significantly deteriorates, whereas M-Pilot maintains stable performance and surpasses PAG by an increasing margin. For LaMP-4, both M-Pilot and PAG exhibit similar trends, but M-Pilot consistently outperforms PAG by a substantial and steady margin. These results demonstrate the efficacy of IDPO in enhancing the summarization capabilities of the black-box controller, especially when dealing with varying and plenty of profiles.
B.2 Experiments on challenging MATH500
Figure 7: Examples of M-Pilot performance on MATH500 dataset.
In this section, we show that M-Pilot can generalize well to other more challenging tasks, MATH500 [17] for instance. As shown in Figure. 7, even with single-turn outer-loop interactions, our method already significantly outperforms the black-box only approach. This is because the controller model provides high-quality intermediate guidance, directing the black-box LLM to adjust its execution steps more effectively toward the correct direction. Moreover, when employing multi-turn outer-loop interactions, performance is further enhanced, demonstrating the controller’s ability to refine its instructions through feedback for more efficient guidance of the black-box LLM.
B.3 Comparison with other baselines
We additionally conduct a comprehensive comparison with several state-of-the-art baselines across various domains to demonstrate the effectiveness and generalizability of our proposed framework. In personalization tasks, ORPO [54] optimizes prompts by iteratively generating and evaluating solutions using LLMs without training, while Fermi [22] personalizes LLMs by iteratively refining user-specific prompts based on profiles and misaligned responses; For reasoning tasks, LEAP [58] improves prompts by learning from previous mistakes, whereas Bbox-Adapter [42] employs Noise Contrastive Estimation(NCE) to train an adapter that guides the policy more effectively.
Table 5: Comparison with baselines on LaMP.
Tasks ($\downarrow$) Methods ($\rightarrow$) Fermi [22] ORPO [54] M-Pilot (Ours)
LaMP-2M $\uparrow$ 37.8 34.3 47.0
LaMP-3 $\downarrow$ 34.0 57.0 29.0
Table 6: Comparison with baselines on GSM8K.
Task ($\downarrow$) Methods ($\rightarrow$) LEAP(low-level) [58] LEAP(high-level) [58] Bbox-Adapter [42] M-Pilot (Ours)
GSM8K 77.4 76.6 74.94 93.1
We show the comparison results in Table 5 and Table 6, where all baseline results are reported as in their original papers, and gpt-3.5-turbo is used consistently across all experiments. Our method outperforms the strongest baseline on the LaMP benchmark by 9.2% and 5.0% for LaMP-2M and LaMP-3, respectively. On the GSM8K dataset, our approach achieves a 15.7% improvement over the best baseline, demonstrating the broad effectiveness and versatility of M-Pilot.
From an algorithmic standpoint, ORPO and LEAP both share the core idea of prompt optimization using LLMs, aligning with our motivation. However, they both rely on training-free, in-context learning approaches. In contrast, BBox-Adapter focuses on constraining the black-box LLM’s tree search process rather than providing high-level task guidance. M-Pilot distinguishes itself by introducing a trainable white-box controller for prompt optimization, offering a more powerful and flexible mechanism to steer the black-box LLM and further enhance its performance.
B.4 M-Pilot with other Iterative Training Method
We want to emphasize that the core contribution of our work lies in adapting the pioneering controller-generator framework. This framework is highly flexible and enables controllable multi-turn generation, significantly enhancing the ability of black-box large language models to handle complex, long-horizon tasks. Although we previously use IDPO to demonstrate our approach, we want to clarify that the specific RL method used to train the controller model is merely a design choice and not the primary focus of our work. To demonstrate the adaptability of our framework, we conducted experiments on all three previously mentioned datasets using SimPo [25] as the RL method. Specifically, given input $x$ and generated guidance pair $(g^{+},g^{-})$, the SimPo loss for optimizing the white-box LLM controller can be expressed as:
$\mathcal{L}_{\text{SimPO}}$ $=\mathbb{E}_{(x,g^{+},g^{-})\sim\mathcal{D}}\Biggl[-\log\sigma\Biggl(\eta^{-1}\Biggl(\frac{1}{|g^{+}|}\log\pi_{\theta}(g^{+}|x)-\frac{1}{|g^{-}|}\log\pi_{\theta}(g^{-}|x)-\gamma\Biggr)\Biggr)\Biggr].$ (12)
Remaining consistent with the previous notation, we can then easily represent the training objective for iterative guidance optimization with the SimPo loss as follows:
$\mathcal{L}_{\text{ISimPO}}$ $=\mathbb{E}_{(x,g^{+},g^{-})\sim\mathcal{D}}\Biggl[-\log\sigma\Biggl(\eta^{-1}\Biggl(\frac{1}{|g^{+}|}\log\pi_{\theta}^{(m+1)}(g^{+}|x)-\frac{1}{|g^{-}|}\log\pi_{\theta}^{(m+1)}(g^{-}|x)-\gamma\Biggr)\Biggr)\Biggr].$ (13)
Table 7: Additional experimental results on the personalization task using the LaMP benchmark. All baseline settings and notations remain consistent with main experiment. ISimPo represents Iterative SimPo loss.
Dataset ($\rightarrow$) LaMP-1 LaMP-2N LaMP-2M LaMP-3 LaMP-4
gpt-4o-mini 0.514 0.513 0.655 0.473 0.413 0.325 0.371 0.673 0.132 0.116 0.992
RAG (k=1) [34] 0.626 0.624 0.733 0.539 0.444 0.378 0.311 0.631 0.141 0.126 1.296
RAG (k=4) [34] 0.632 0.632 0.792 0.611 0.502 0.430 0.272 0.579 0.161 0.146 2.953
PAG [32] 0.624 0.624 0.775 0.559 0.496 0.443 0.316 0.645 0.143 0.130 1.968
M-Pilot (IDPO) 0.640 0.639 0.823 0.607 0.527 0.465 0.277 0.581 0.174 0.160 4.298
M-Pilot (ISimPo) 0.628 0.628 0.826 0.598 0.522 0.461 0.294 0.614 0.180 0.167 4.997
Table 8: Additional experimental results on GSM8K dataset.
Dataset ($\rightarrow$) GSM8K GSM-HARD
Method ($\downarrow$) gpt-3.5 4o-mini gpt-3.5 4o-mini
CoT 0.809 0.932 0.406 0.500
Least-to-Most 0.811 0.908 0.425 0.498
PAL 0.802 0.920 0.638 0.748